Project

General

Profile

Bug #1876

Muon Beamline Step IV prep list

Added by Nebrensky, Henry over 7 years ago. Updated over 7 years ago.

Status:
Closed
Priority:
Normal
Category:
Beam Line
Start date:
21 September 2016
Due date:
07 November 2016
% Done:

100%

Estimated time:

Description

Muon Beamline to-do list


Related issues

Follows Operations - Feature #1854: Muon Beamline Step IV prep listClosedNebrensky, Henry06 June 201628 June 2016

Actions
Precedes Operations - Bug #1892: Muon Beamline Step IV prep listClosedNebrensky, Henry08 November 201615 February 2017

Actions
#1

Updated by Nebrensky, Henry over 7 years ago

  • Description updated (diff)
#2

Updated by Nebrensky, Henry over 7 years ago

Stuff that needs sorting out.
Important stuff in bold

Target

  • Light Loss (TGT) - stabilised in mild saturation again
  • Trace fibre connections - January shutdown awaiting MICE Note
  • Return/replace target1ctl
  • S3 and T2 - assembly of next spare (TGT) - pinch point is getting optics block off T1
  • Clean-up of R78 target control crate (DL) need to restart VME control (and possible move to R9)
  • ISIS Target - cooling water flow read-out?
  • Target Alarm handler #1613 update 22 (CAM)
  • ISIS data items #1669 (CAM)
  • Telephone link #977
  • Activation Study (activation seen: limit of 2 V at MS/64 stands)
  • Beambump tune
  • Updated target manual (EO/JL)

Conventional Magnets

  • COOLING WATER
    • sufficient flow to heat exchanger ( #1875 )
    • complete new demin water circuit ( #1878 )
  • review existing water flows and temperature range (MBC)
    • need max current limits for quads in control software? (CAM)
  • Calibrate D1 & D2 current reading and update IOC software (DL)
  • Q4 Magnet-On lamp and general replacement - Andy Gallagher
  • 50 A minimum current limits for D1, D2 in control software (CAM)
  • Diagnose flooded Danfysik 9000 supply
    • faulty PCB - Richard Hale contacting Danfysik about a replacement
    • Have ordered replacement board
    • New board tested and base board reworked
    • Looking for suitable storage space
  • Q1 alarm on -ve voltage at zero amps (CAM)
    • Now on both Q1 and Q2 MPS
  • Q2 start-up recalcitrance
  • Q4 Earth Leakage (MBC) - D2,Q4,Q5 flushed through
    • Q4 threshold loosened
    • Repeat Earth Leakage measurements next shutdown
  • Q5 leaks fitting tightened, the other has disappeared
  • Q6 leak leave until time
  • Leak sensors and drip trays
  • Protection for D2 klixons and cabling
  • "All Off" controls misbehaviour
  • Water, temperatures and MPS water flows/temperatures in ALH/archiver (CAM) #1880
    • Interlocks alarm is inoperative
  • Update magnet and POMPOM user guides (MBC)
  • Annual POMPOM radiation damage test (MBC)
    • negative
    • positive

Decay Solenoid

  • LHe level sensor
  • Replace shaft bearing, again...
  • Overheating investigation (DK/DL)
  • Replace PSU (DK/DL)
    • Smoothing of read-back current
    • Calibrate current reading and update IOC software (DL)
  • 120A trip investigation (DK/DL)
  • Replace final battery cell (DL)
  • Diodes in BBU circuit (DL) - for now, read-back changed to avoid false alarms
  • Boosted air (MBC)
    • Trench sensor (awaiting read-out)
    • Compressor sensor
    • Backup compressor
  • Linde PC install: awaiting validation and IOC (DK) - licence issues #1864
    • MLCR monitor and KVM implications (HN)

Proton Absorber

  • Combine survey documents (MBC)

Beamstop

  • Translate Survey document (MBC)

Diffuser

  • Air crate: solenoid valve (Oxf)
    • Failed PROM in electronics crate; back in Oxford

General

  • BLOC training and experience (MBC)
  • Shared services/environment (water, air, aircon, Hall temp., computing/networking, UPS ( #1213 ) in Alarm Handler and in Archiver (CAM) #1881

Beyond the Muon Beamline

#3

Updated by Nebrensky, Henry over 7 years ago

  • Due date set to 31 October 2016
  • Category set to Beam Line
  • Assignee set to Nebrensky, Henry
#4

Updated by Nebrensky, Henry over 7 years ago

  • Due date changed from 31 October 2016 to 07 November 2016
#5

Updated by Nebrensky, Henry over 7 years ago

  • Status changed from Open to Closed
  • % Done changed from 0 to 100

Follow up in #1892

Also available in: Atom PDF